The length of a rectangle is 1 inch. less than twice the width. the area of the rectangle is 21 spuare inches. What is the length of the rectangle.
Answer by checkley77(12844)

L=2W-1
W(2W-1)=21
2W^2-W-21=0
(2W-7)(W+3)=0
2W-7=0
2W=7
X=7/2
X=3.5 FOR THE WIDTH.
L=2*3.5-1
L=7-1
L=6 FOR THE LENGTH.
6*3.5=21
21=21